# CulturePark: Boosting Cross-cultural Understanding in Large Language Models

### v1

@@ Abstract
Cultural bias is pervasive in many large language models (LLMs), largely due to the deficiency of data representative of different cultures. Typically, cultural datasets and benchmarks are constructed either by extracting subsets of existing datasets or by aggregating from platforms such as Wikipedia and social media. However, these approaches are highly dependent on real-world data and human annotations, making them costly and difficult to scale. Inspired by cognitive theories on social communication, this paper introduces CulturePark, an LLM-powered multi-agent communication framework for cultural data collection. CulturePark simulates cross-cultural human communication with LLM-based agents playing roles in different cultures. It generates high-quality cross-cultural dialogues encapsulating human beliefs, norms, and customs. Using CulturePark, we generated 41,000 cultural samples to fine-tune eight culture-specific LLMs. We evaluated these models across three downstream tasks: content moderation, cultural alignment, and cultural education. Results show that for content moderation, our GPT-3.5-based models either match or outperform GPT-4 on $41$ datasets. Regarding cultural alignment, our models surpass GPT-4 on Hofstede’s VSM 13 framework [1] . Furthermore, for cultural education of human participants, our models demonstrates superior outcomes in both learning efficacy and user experience compared to GPT-4. CulturePark proves an important step in addressing cultural bias and advancing the democratization of AI, highlighting the critical role of culturally inclusive data in model training.
@@ 1 Introduction
Culture is an important part of human society, composed of human beliefs, norms, customs etc. [2]. As large language models (LLMs) play a vital role in daily communication, recommender systems [3, 4], and education [5, 6], it is imperative for LLMs to perceive and reflect different cultures. However, current state-of-the-art LLMs are reported to be biased towards mainstream culture while ignoring others, resulting in a cultural bias problem [7, 8, 9, 10, 11, 12]. This leads to stereotypical impressions of different cultures, which can even exacerbate social conflicts [13]. The main reason behind cultural bias is that the training corpus of LLMs is dominated by English data that expresses the cultural values and opinions of Western people. Much less can be learned about other cultures simply because there is less data available, i.e., a low-resource situation.
Existing approaches to solve the cultural bias problem in LLMs include prompt engineering [14, 11, 15] and pre-training on non-English languages [16, 17, 18, 19, 20, 21]. Prompt engineering consists in tuning prompts for different cultural tasks, but the benefits do not hold reliably across various downstream tasks. Pre-training in various languages is promising but the data collection and the pre-training itself are both very costly. More importantly, cultural differences are embodied in many aspects such as opinions, customs, norms, and languages. One model serving all cultures may meet cultural conflict and misalignment problems [7, 8, 9]. Thus, it is necessary to fine-tune culture-specific models that target specific cultures. Recently, Li et al. [22] proposed CultureLLM, which augments the fine-tuning data of LLMs via semantic data augmentation to train culture-specific LLMs. However, the generated data lacks diversity because it is implemented by generating semantically equivalent sentences of seed examples.
@@
From the perspective of cognitive social science, our framework is inspired by the Cognitive Conflict Theory (CCT) [23, 24] and Social Cognitive Theory (SCT) [25] to foster a collaborative and communicative environment for mutual understanding of cultures. Specifically, CulturePark allows agents to encounter cognitive conflicts, which trigger deeper thinking on certain topics according to CCT [23, 24]. At the same time, a deeper understanding of cultures can be prompted through interaction and communication with other agents, as suggested by SCT [25]. In favor of these theories, we found that CulturePark triggers LLMs’ cross-cultural understanding ability, boosts novel opinions by allowing agents to think deeper, and benefits data augmentation by creating more comprehensive answers to the questions. CulturePark has the potential to help cultural data collection, cultural alignment, and improve AI democracy.
@@
    Leveraging CulturePark, we generate and augment novel questions and more comprehensive answers, leading to $41K$ cultural samples in total. Those data contain rich and diverse information on different aspects of culture such as norms, opinions, and backgrounds. We then fine-tune cultural specific LLMs for different cultures.
@@
    We evaluate CulturePark in three key experiments: 1) The fine-tuned LLMs outperforms GPT-4 in $5$ cultures on $26$ content moderation tasks and approach GPT-4 on other tasks; 2) Our fine-tuned LLMs achieve better performance on cultural alignment experiments via Hofstede’s cultural dimensions theory [26]; and 3) Human participants can perform more effective culture learning in a situated learning experiments and show better satisfaction compared to GPT-4.
@@ 2.1 Cultural bias in LLMs
A body of research has explored cultural biases in LLMs. Johnson et al. [27] examined conflicts in model outputs and input values, using moral value pluralism to analyze GPT-3’s responses against global demographics. Their results showed that conflicting values were more aligned with reported dominant US values. [10] highlighted a bias towards Western culture in models processing Arabic, exacerbated by English-aligned prompts, suggesting mitigation through cultural tokens. The Cultural Alignment Test (CAT), based on Hofstede’s framework [26], evaluated cultural values in models like ChatGPT and Bard across different cultures, revealing the highest cultural alignment for GPT-4 with US values [9]. [8] found that ChatGPT aligned well with American culture but struggled with other cultures, particularly under English prompts. Additionally, [7] reported that multilingual LLMs had limited abilities to reason with proverbs and exhibited a “culture gap” in handling translations, leading to the development of the MAPS dataset for assessing proverb comprehension in six languages.
@@ 2.2 Cultural benchmarks and datasets
Extensive research has focused on developing cultural benchmarks, which can be categorized into two types: collecting existing datasets and synthesizing new ones. First, most work adopted existing datasets as sources of cultural data. [11] introduced a benchmark utilizing cultural items to analyze cultural dominance, based on sources like WVS [28] and PCT [29]. Later work includes Cultural Alignment Test [9], NORMSAGE [30], WorldValueBench [31], and NORMAD [32] that sourced from different existing datasets. Other types of data sources include CultureAtlas [33] and MAPS [7] which collected data from Wikimedia; Candle [34] and CultureBank [35] sourced their data from social media such as Tiktok and Reddit. In contrast, there was an emerging trend to perform data augmentation for cultural LLMs. Li et al. [22] proposed semantic data augmentation to synthesize cultural data by enriching the semantic equivalence of the generated samples.
CulturePark significantly differs from those that perform direct data collection from existing datasets; it also differs from CultureLLM [22] since CulturePark leverages multi-agent communication for data generation, which is more natural and can generate more diverse datasets.
@@ 2.3 Existing solutions to cultural bias
There are primarily two types of approaches to addressing the problem of cultural bias: prompt engineering and pre-training. The work of [14, 11] viewed LLMs as amalgamations of cultural perspectives, which can be steered towards specific cultural viewpoints through prompt engineering. In contrast, Rao et al. [15] integrated cultural values directly into the prompts. Although prompt engineering is cost-effective, its efficacy is questionable, particularly in low-resource cultures where LLMs may lack relevant cultural knowledge due to underrepresentation in pre-training data. An alternative strand of research focuses on pre-training and fine-tuning [16, 17, 18, 19, 20, 21]. Those approaches developed culturally aware LLMs for various cultures by assembling large-scale pre-training datasets, followed by fine-tuning to enhance alignment. Despite achieving significant performance improvements, those methods were both costly and time-consuming, making them impractical for broader application across numerous cultures and countries. Furthermore, they still face challenges in low-resource cultures where acquiring pre-training data is difficult. For instance, the MaLA-500 [36] aimed to train a new LLM on Llama 2 to support $534$ languages, illustrating the resource-intensive nature of this approach. Different from those approaches, CulturePark provides a cost-efficient solution to the cultural bias problem which includes data augmentation and fine-tuning.
@@ 3.1 Design
CulturePark is an LLM-powered [^4] We use GPT-3.5-Turbo in this work due to its great performance, high efficiency, and manageable price. CulturePark is a flexible platform that naturally supports other LLMs such as GPT-4 and Llama models. More importantly, CulturePark allows to use different LLMs for the main contact and delegate (e.g., GPT-3.5 for main contact and Llama-2 for delegate), which makes it flexible to extend to other models and evaluate the cross-model understanding ability. cross-cultural communication framework that generates data to support culture-related research such as building cultural-specific LLMs and performing cultural alignment. It is inspired by Cognitive Conflict Theory (CCT) and Social Cognition Theory (SCT) to design multi-turn communications for a deeper understanding of cultural topics. CCT posits that cognitive conflicts can help individuals engage more in deeper thinking [23, 24], and SST emphasizes that individuals can deepen their understanding of perspectives through explanation and debate [25].
@@ 3.2 Data refinement and fine-tuning
The seed questions initiating the communication have two sources: World Values Survey (WVS) [28] and Global Attitudes surveys (GAS) from Pew Research Center [37]. WVS is a global research project that explores people’s beliefs and values worldwide, examining how these beliefs evolve over time. Pew Research Center, a nonpartisan organization, provides data and research on public opinion, social issues, and demographic trends both in the U.S. and globally. Its Global Attitudes surveys cover a wide range of topics, including politics, media, technology, religion, race, and ethnicity. In total, we select $4.1$k seed data and generate $41$k dialogues (each dialogue contains several sentences). We show the details on data numbers for different cultures in Table 8. We also conduct a statistics analysis on the dataset leveraging GPT-4. As summarized in Figure 8, the dataset contains human belief (59.68%), norm (29.54%) and custom (10.78%) involving data on $8$ different cultures. Figures 7 and 10(a) show some examples of the seed data and the generated dialogues.
The generated dataset may not be directly used for fine-tuning since it could contain redundant and incorrect information that should be handled. As shown in Figure 2(b), we design data refinement to refine the dataset. First, the opinions on target culture are extracted from the generated dialogues via GPT-4, such as “The Arabian equates their parents’ happiness and satisfaction to their own success” and “The Arabian emphasize Sabr, which is about showing resilience, maintaining a positive attitude and having faith during difficult times”. Second, several extracted opinions could be irrelevant to the initial question or contradict with seed data, motivating us to perform verification to reserve only highly related opinions. Furthermore, since the generated data could be semantically similar, we remove the redundant samples to improve the diversity. To be specific, we get sentence embeddings via text-embedding-3-small [38] and cluster the embedding using K-means. We reserve one sample for each cluster as representative data. Eventually, we get the high-quality cultural data for different cultures. The ablation of refinement is in Table 2.
Algorithm 1 shows the pipeline for data refinement. After refinement, there are $41$k samples (input-output pairs) left for fine-tuning, i.e., one sample for one dialogue. Examples of the samples are provided in Figure 11. Afterwards, we can fine-tune cultural-specific LLMs using either open-source LLMs or fine-tuning service. In this paper, we mainly use OpenAI API to fine-tune GPT-3.5-Turbo due to its efficiency. The hyperparameters are shown in Table 6. We further provide fine-tuning experiments on Llama2-70b in Section 5.4.
(a) Cross-cultural dialogue dataset
(b) Data refinement for fine-tuning
Figure 2: Cross-cultural dialogue and data refinement for fine-tuning LLMs using CulturePark.
@@ 3.3 CulturePark benefits cultural understanding and fine-tuning
Communication triggers LLMs’ cross-cultural understanding ability. We observed that agents try to understand each other’s opinions and the reasons sourced from their different cultural backgrounds. For example, the blue sentences in Figure 2(a) show cross-cultural understanding ability of LLMs, such as “I do understand and respect the sentiment of wanting to make parents proud, as they have often sacrificed a lot for their children”. Leveraging GPT-4-Turbo, we analyzed the topics in the dataset such as human beliefs, norms, and customs, which can be further used as data collections for building culturally specific models. Section A.2 shows the details of the dialogue dataset, indicating that the generated topics are mostly about culture. Then, we randomly sampled $750$ dialogues for each culture and evaluated the communication using prompts in Appendix D. As summarized in Table 4, on average, the ratio of statements that express cross-cultural understanding is $80.80$%. The analysis also verifies the effectiveness of CulturePark in extending topics and cross-cultural understanding.
@@
CulturePark naturally assists cultural data augmentation by creating novel questions and comprehensive answers. On the one hand, agents in different cultures can generate new opinions towards certain topics, which intuitively diversifies the input questions. On the other hand, the initial seed data only contains short answers such as “Strongly agree” with no further explanations. Our platform allows deeper and more comprehensive communication of agents, thus generating more detailed answers such as “Strongly agree. I believe that pleasing parents and elders is a sign of respect and love” and “Strongly agree. I equate my parents’ happiness and satisfaction to my own success”. Additionally, agents can extend the topics that conflict with their own opinions and provide more informative evidence to support their viewpoints. This strategy helps to generate informative and diverse data continuously. Section 5.1 presents some detailed results on diversity gain, showing that the generated data has significantly larger diversity.

### v2

@@ Abstract
Cultural bias is pervasive in many large language models (LLMs), largely due to the deficiency of data representative of different cultures. Typically, cultural datasets and benchmarks are constructed either by extracting subsets of existing datasets or by aggregating from platforms such as Wikipedia and social media. However, these approaches are highly dependent on real-world data and human annotations, making them costly and difficult to scale. Inspired by cognitive theories on social communication, this paper introduces CulturePark, an LLM-powered multi-agent communication framework for cultural data collection. CulturePark simulates cross-cultural human communication with LLM-based agents playing roles in different cultures. It generates high-quality cross-cultural dialogues encapsulating human beliefs, norms, and customs. Using CulturePark, we generated 41,000 cultural samples to fine-tune eight culture-specific LLMs. We evaluated these models across three downstream tasks: content moderation, cultural alignment, and cultural education. Results show that for content moderation, our GPT-3.5-based models either match or outperform GPT-4 on $41$ datasets. Regarding cultural alignment, our models surpass GPT-4 on Hofstede’s VSM 13 framework (Hofstede, 2013) . Furthermore, for cultural education of human participants, our models demonstrate superior outcomes in both learning efficacy and user experience compared to GPT-4. CulturePark proves an important step in addressing cultural bias and advancing the democratization of AI, highlighting the critical role of culturally inclusive data in model training. Code is released at https://github.com/Scarelette/CulturePark.
@@ 1 Introduction
Culture is an important part of human society, composed of human beliefs, norms, customs, etc. (Spencer-Oatey and Franklin, 2012). As large language models (LLMs) play a vital role in daily communication (Yang et al., 2024), recommendation systems (Li et al., 2023; Fan et al., 2023), and education (Shaikh et al., 2023), it is imperative for LLMs to perceive and reflect different cultures. However, current state-of-the-art LLMs have been reported to be biased towards mainstream culture while ignoring others, resulting in a cultural bias problem (Liu et al., 2023; Cao et al., 2023; Masoud et al., 2023; Naous et al., 2023; Wang et al., 2023; Johnson et al., 2022b). This leads to stereotypical impressions of different cultures, which can even exacerbate social conflicts (Ryan et al., 2024). The main reason behind cultural bias is that the training corpus of LLMs is dominated by English data that express the cultural values and opinions of western people. Much less can be learned about other cultures simply because there is less data available, i.e., a low-resource situation.
Existing approaches to solve the cultural bias problem in LLMs include prompt engineering (Kovač et al., 2023; Wang et al., 2023; Rao et al., 2023) and pre-training in non-English languages (Pires et al., 2023; Chan et al., 2023; Nguyen et al., 2023b; Pipatanakul et al., 2023; Abbasi et al., 2023; Lin and Chen, 2023). Prompt engineering consists of tuning prompts for different cultural tasks, but the benefits do not hold reliably across various downstream tasks. Pre-training in various languages is promising, but the data collection and the pre-training itself are both very costly. More importantly, cultural differences are embodied in many aspects such as opinions, customs, norms, and languages. A model serving all cultures may face cultural conflict and misalignment problems (Liu et al., 2023; Cao et al., 2023; Masoud et al., 2023). Thus, it is necessary to fine-tune culture-specific models that target specific cultures. Recently, Li et al. (2024) proposed CultureLLM, which augments the fine-tuning data of LLMs via semantic data augmentation to train culture-specific LLMs. However, the generated data lack diversity because it is implemented by generating semantically equivalent sentences of seed examples.
@@
From the perspective of cognitive social science, our framework is inspired by Cognitive Conflict Theory (CCT) (Limón, 2001; Cosier and Rose, 1977) and Social Cognitive Theory (SCT) (Fiske and Taylor, 1991) to foster a collaborative and communicative environment for mutual understanding of cultures. Specifically, CulturePark allows agents to encounter cognitive conflicts, which triggers deeper thinking on certain topics according to CCT (Limón, 2001; Cosier and Rose, 1977). At the same time, a deeper understanding of cultures can be prompted through interaction and communication with other agents, as suggested by SCT (Fiske and Taylor, 1991). In favor of these theories, we found that CulturePark triggers LLMs’ cross-cultural understanding ability, boosts novel opinions by allowing agents to think deeper, and benefits data augmentation by creating more comprehensive answers to the questions. CulturePark has the potential to facilitate the data collection of culture-related tasks, cultural value alignment, and improving AI democracy.
@@
    Leveraging CulturePark, we generate and augment novel questions and more comprehensive answers, leading to $41K$ cultural samples in total. Those data contain rich and diverse information on different aspects of culture, such as norms, opinions, and backgrounds. We then fine-tune cultural specific LLMs for different cultures.
@@
    We evaluate CulturePark in three key experiments: 1) The fine-tuned LLMs outperforms GPT-4 in $5$ cultures on $26$ content moderation tasks and approach GPT-4 on other tasks; 2) Our fine-tuned LLMs achieve better performance on cultural alignment experiments via Hofstede’s cultural dimensions theory (Geert Hofstede, 2010); and 3) Human participants can perform more effective culture learning in situated learning experiments and show better satisfaction compared to GPT-4.
@@ 2.1 Cultural bias in LLMs
A body of research has explored cultural biases in LLMs. Johnson et al. (2022a) examined conflicts in model outputs and input values, using moral value pluralism to analyze the responses of GPT-3 against global demographics. Their results showed that conflicting values were more aligned with the dominant US values reported. Naous et al. (2023) highlighted a bias towards Western culture in models processing Arabic, exacerbated by English-aligned prompts, suggesting mitigation through cultural tokens. The Cultural Alignment Test (CAT), based on Hofstede’s framework (Geert Hofstede, 2010), evaluated cultural values in models such as ChatGPT and Bard across different cultures, revealing the highest cultural alignment for GPT-4 with US values (Masoud et al., 2023). Cao et al. (2023) found that ChatGPT aligned well with American culture but struggled with other cultures, particularly under English prompts. Additionally, Liu et al. (2023) reported that multilingual LLMs had limited abilities to reason with proverbs and exhibited a “culture gap” in handling translations, leading to the development of the MAPS dataset for assessing proverb comprehension in six languages.
@@ 2.2 Cultural benchmarks and datasets
Extensive research has focused on developing cultural benchmarks, which can be categorized into two types: collecting existing datasets and synthesizing new ones. First, most of the work adopted existing datasets as sources of cultural data. Wang et al. (2023) introduced a benchmark that uses cultural items to analyze cultural dominance, based on sources such as WVS (Survey, 2022b) and PCT (Mudde, 2016). Later work includes Cultural Alignment Test (Masoud et al., 2023), NORMSAGE (Fung et al., 2022), WorldValueBench (Zhao et al., 2024), and NORMAD (Rao et al., 2024) that sourced from different existing datasets. Other types of data sources include CultureAtlas (Fung et al., 2024) and MAPS (Liu et al., 2023) which collected data from Wikimedia; Candle (Nguyen et al., 2023a) and CultureBank (Shi et al., 2024) sourced their data from social media such as Tiktok and Reddit. In contrast, there was an emerging trend to perform data augmentation for cultural LLMs. Li et al. (2024) proposed semantic data augmentation to synthesize cultural data by enriching the semantic equivalence of the generated samples.
CulturePark significantly differs from those that perform direct data collection from existing datasets; it also differs from CultureLLM (Li et al., 2024) since CulturePark leverages multi-agent communication for data generation, which is more natural and can generate more diverse datasets.
@@ 2.3 Existing solutions to cultural bias
There are primarily two types of approach to addressing the problem of cultural bias: prompt engineering and pre-training. The work of (Kovač et al., 2023; Wang et al., 2023) viewed LLMs as amalgamations of cultural perspectives, which can be oriented toward specific cultural perspectives through prompt engineering. In contrast, Rao et al. (Rao et al., 2023) integrated cultural values directly into the prompts. Although prompt engineering is cost-effective, its efficacy is questionable, particularly in low-resource cultures where LLMs may lack relevant cultural knowledge due to underrepresentation in pre-training data. An alternative strand of research focuses on pre-training and fine-tuning (Pires et al., 2023; Chan et al., 2023; Nguyen et al., 2023b; Pipatanakul et al., 2023; Abbasi et al., 2023; Lin and Chen, 2023). Those approaches developed culturally aware LLMs for various cultures by assembling large-scale pre-training datasets, followed by fine-tuning to enhance alignment. Despite achieving significant performance improvements, these methods were both costly and time consuming, making them impractical for a broader application across numerous cultures and countries. Furthermore, they still face challenges in low-resource cultures where acquiring pre-training data is difficult. For example, MaLA-500 (Lin et al., 2024) aimed to train a new LLM in Llama 2 to support $534$ languages, illustrating the resource-intensive nature of this approach. Unlike these approaches, CulturePark provides a cost-effective solution to the cultural bias problem, including data augmentation and fine-tuning.
@@ 3.1 Design
CulturePark is an LLM-powered [^2] We use GPT-3.5-Turbo in this work due to its great performance, high efficiency, and manageable price. CulturePark is a flexible platform that naturally supports other LLMs such as GPT-4 and Llama models. More importantly, CulturePark allows to use different LLMs for the main contact and delegate (e.g., GPT-3.5 for main contact and Llama-2 for delegate), which makes it flexible to extend to other models and evaluate the cross-model understanding ability. cross-cultural communication framework that generates data to support culture-related research such as building cultural-specific LLMs and performing cultural alignment. It is inspired by Cognitive Conflict Theory (CCT) and Social Cognition Theory (SCT) to design multi-turn communications for a deeper understanding of cultural topics. CCT posits that cognitive conflicts can help individuals engage more in deeper thinking (Limón, 2001; Cosier and Rose, 1977), and SST emphasizes that individuals can deepen their understanding of perspectives through explanation and debate (Fiske and Taylor, 1991).
@@ 3.2 Data refinement and fine-tuning
The seed questions initiating the communication have two sources: World Values Survey (WVS) (Survey, 2022b) and Global Attitudes surveys (GAS) from Pew Research Center (Survey, 2022a). WVS is a global research project that explores people’s beliefs and values worldwide, examining how these beliefs evolve over time. Pew Research Center, a nonpartisan organization, provides data and research on public opinion, social issues, and demographic trends both in the U.S. and globally. Its Global Attitudes surveys cover a wide range of topics, including politics, media, technology, religion, race, and ethnicity. In total, we select $4.1$k seed data and generate $41$k dialogues (each dialogue contains several sentences). We show the details of the data numbers for different cultures in Table 8. We also performed a statistical analysis on the GPT-4-based dataset. As summarized in Figure 10, the dataset contains human belief (59.68%), norm (29.54%) and custom (10.78%) involving data on $8$ different cultures. Figures 9 and LABEL:fig-case-control show some examples of the seed data and the generated dialogues.
The generated dataset may not be directly used for fine-tuning since it could contain redundant and incorrect information that should be handled. As shown in LABEL:fig-pipeline-2, we design data refinement to refine the dataset. First, the opinions on target culture are extracted from the dialogues generated via GPT-4, such as “The Arabian equates their parents’ happiness and satisfaction to their own success” and “The Arabian emphasize Sabr, which is about showing resilience, maintaining a positive attitude and having faith during difficult times”. Second, several extracted opinions could be irrelevant to the initial question or contradict with seed data, motivating us to perform verification to reserve only highly related opinions. Furthermore, since the generated data could be semantically similar, we remove redundant samples to improve diversity. To be specific, we obtain sentence embeddings via text-embedding-3-small (OpenAI, 2024) and cluster the embedding using K-means. We reserve one sample for each cluster as representative data. Eventually, we get the high-quality cultural data for different cultures. The ablation of the refinement is in Table 2.
Algorithm 11 shows the pipeline for data refinement. After refinement, there are $41$k samples (input-output pairs) left for fine-tuning, i.e., one sample for one dialogue. Examples of the samples are provided in Figure 14. Afterwards, we can fine-tune cultural-specific LLMs using either open-source LLMs or fine-tuning service. In this paper, we mainly use OpenAI API to fine-tune GPT-3.5-Turbo due to its efficiency. Hyperparameters are shown in Table 6. We further provide fine-tuning experiments on Llama2-70b in Section 5.4.
@@ 3.3 CulturePark benefits cultural understanding and fine-tuning
Communication triggers LLMs’ cross-cultural understanding ability. We observed that agents try to understand each other’s opinions and the reasons sourced from their different cultural backgrounds. For example, the blue sentences in LABEL:fig-pipeline-1 show cross-cultural understanding ability of LLMs, such as “I do understand and respect the sentiment of wanting to make parents proud, as they have often sacrificed a lot for their children”. Leveraging GPT-4-Turbo, we analyzed the topics in the dataset such as human beliefs, norms, and customs, which can be further used as data collections for building culturally specific models. Section B.2 shows the details of the dialogue dataset, indicating that the generated topics are mostly about culture. Then, we randomly sampled $750$ dialogues for each culture and evaluated the communication using the prompts in Appendix E. As summarized in Table 4, on average, the ratio of statements that express cross-cultural understanding is $80.80$%. The analysis also verifies the effectiveness of CulturePark in expanding topics and cross-cultural understanding.
@@
CulturePark naturally assists cultural data augmentation by creating novel questions and comprehensive answers. On the one hand, agents in different cultures can generate new opinions towards certain topics, which intuitively diversifies the input questions. On the other hand, the initial seed data only contain short answers such as “Strongly agree” with no further explanations. Our platform allows deeper and more comprehensive communication of agents, thus generating more detailed responses such as “Strongly agree. I believe that pleasing parents and elders is a sign of respect and love” and “Strongly agree. I equate my parents’ happiness and satisfaction to my own success”. Additionally, agents can extend the topics that conflict with their own opinions and provide more informative evidence to support their viewpoints. This strategy helps to generate informative and diverse data continuously. Section 5.1 presents some detailed results on diversity gain, showing that the generated data has significantly larger diversity.